OpenStudy (technodynamic):

Quick question: My final exam today is worth 25% of my final grade, what score would I need to get on the final to pass with a 4.0??? So far I have: 95/100 on Exam 1. 90/100 on Exam 2.

OpenStudy (technodynamic):

I'm really taking a final today.

OpenStudy (lgbasallote):

it depends on the range of 4.0 is it 95-100?

OpenStudy (technodynamic):

4.0 ranges from 90-100

OpenStudy (anonymous):

You have to get at least the lowest range of an A

OpenStudy (lgbasallote):

it actually depends on the percentage of the final exam...percentage of the quizzes..and all other rubrics...

OpenStudy (anonymous):

You have two other exams that weight more than your final exam?

OpenStudy (lgbasallote):

so as for your problem i'll choose E. not enough information given :P

OpenStudy (anonymous):

How is the grade determined? Total points obtained / Total points available? Or 75% based on exams and 25% based on final?

OpenStudy (technodynamic):

75% for exams one and two. 25% for the final.

OpenStudy (lgbasallote):

average of 2 quizzes * 0.75 + 0.25x = 90 ave = 185/2 = 92.5 92.5 * 0.75 + 0.25x = 90 0.25x = 90 - 69.375 0.25x = 20.625 x = 82.5 pts

OpenStudy (lgbasallote):

these are the times i love algebra \m/ hehe

OpenStudy (anonymous):

OK, and does the class depend upon points earned? e.g. you earned 95 points of 100 on exam1, so your average is a 95, then you earned 90 out 100 on exam2, so you earned a total of 185/200 = .925. If we assume it just works off the final grade, then you have the equation: (.75)(185/200) + (.25)(FINAL EXAM points earned/points total) = .9 (.75) * (.925) + (.25) * (FPE/FPT) = .9 (.9 - (.75)*(.925))/.25 = (FPE/FPT)

OpenStudy (technodynamic):

Yes points earned, and Thanks guys!

OpenStudy (anonymous):

But that you can aim for higher than .9, in fact aiming for a 100 would be the best scenario of course ;)

OpenStudy (technodynamic):

lol yeah for sure. I hope to pass the final lol

OpenStudy (inkyvoyd):

Are averages weighted? or are they all the same?

OpenStudy (inkyvoyd):

If one test counts more than another, then you have to do calculations differently.

OpenStudy (lgbasallote):

lol i think you pass the subject without the finals already :P what subject is that?

OpenStudy (inkyvoyd):

Nevermind, read the posts above, and your problem is solved.

OpenStudy (technodynamic):

Biochemistry :X lol

OpenStudy (inkyvoyd):

Eh. Biochem doesn't need that much math.

OpenStudy (technodynamic):

Yeah. It's a lot of pressure trying to get decent grades. I'm taking other classes, and the workload they give is almost super tiresome. I think i might do okay.......Hopefully lol.

OpenStudy (inkyvoyd):

Just don't do anything engineering. I mean, that would suck.

OpenStudy (technodynamic):

Engineering requires a lot of math <--my weakest subject lol. I'll stick to science for now. haha